1) Assume that a 3-year treasury security yields 4.10%. Also assume that the real risk-free rate...

1)
Assume that a 3-year treasury security yields 4.10%. Also assume that the real risk-free rate (r*) is 0.75% and inflation is expected to be 2.25% annually for the next 3 years. In addition to inflation, the nominal insterest rate includes a maturity risk premium (MRP) that reflects interest rate risk. What is the maturity risk premium for the 3-year security? Round answer to two decimal places.

2) a treasury bond that matures in 10 years has a yield of 4.75%. A 10 year corporate bond has a yield of 6.50%. Assume that the liquor premium on the corporate bond is 0.6%. What is the default risk premium on the corporate bond? Round answer to two decimal places.
Answer #1

Given in the question:

yield on 3-year treasury security , r = 4.10%

This is the quoted/nominal interest rate

real risk-free rate (r*) = 0.75%

expected inflation for 3 years = 2.25% per year

Now, we know that

Quoted interest rate = Real risk-free rate(r*) + inflation premium (IP) + Maturity risk premium (MRP)

Inflation premium is the average expected premium over the life of the treasury security

since annual expected inflation over the 3 years remains 2.25% per year, the average expected premium over 3 years will also be 2.25% ( ( 2.25% + 2.25% + 2.25%)/3 =2.25%)

Thus inflation premium = 2.25%

Now putting the known values in the above equation

4.10 = 0.75 + 2.25 + MRP

MRP = 4.1-0.75-2.25 = 1.1%